Break-Even Sales and Sales Mix for a Service Compa
Answer #1

Fixed costs

Fuel = 7783

Air crew salaries = 5962

Airplane depreciation = 2815

Total fixed= 16560

Contribution = Sale price - variable cost

Business class = 545 - 45,= 500

Ecomomy class = 255 - 35,= 220

Sales mix = 20 % business class and 80 % economy class

Now, weighted average contribution margin per seat

Business class = Contribution of business class x sales mix

= 500 x 20 %,= 100

Economy class = Contribution of Economy class x sales mix

= 220 x 80 %, =176

Combined weighted contribution per seat = 100 +176 , =276

a) Now seats at breakeven can be calculated as = Fixed cost / combined weighted contribution per seat

= 16560 / 276, = 60 seats

b) Seats in business class = 60 seats x 20 %, =12 seats

economy class =60 seats x 80 %, =48 seats